Grade 12th passMechanics

If the initial velocity of the particle is 2m/s and it attains a velocity of 8m/s in 3s what will be it`s displacement.

Aditya Pathak
8 Years agoGrade 12th pass

1 Answer

DIWAKAR PANDEY
ApprovedApproved Tutor Answer8 Years ago
We will 1st eq. Of motion V=u+at 8=2+a×3 6=3a a=2m/s^2 Now use 2nd eq of motion S=ut+1/2at×t S= 2×3+1/2×2×3×3 S=6+9S=15m